OpenStudy (madinaxxx):

In a closed system, Glider A with a mass of 0.40 kg and a speed of 10.00 m/s collides with Glider B at rest with a mass of 1.20 kg. The two interlock and move together. What speed are they moving? 10.0 m/s 0.5 m/s 2.5 m/s 1.0 m/s

OpenStudy (fortytherapper):

[Physics] First we need to figure out the collision type. If they interlock and move together, is it a elastic or inelastic collision?

OpenStudy (madinaxxx):

inelastic @FortyTheRapper

OpenStudy (fortytherapper):

Right: Since momentum is conserved and they didn't move at a certain angle, we can use \[(m _{1}v _{1})+(m _{2}v _{2}) = (m _{1}+m _{2})v _{f}\]

OpenStudy (madinaxxx):

okay. how do I plug that in?

OpenStudy (fortytherapper):

Let: \[m _{1}=GliderA\] \[m_{2} = GliderB\]

OpenStudy (madinaxxx):

(.40*10)+(1.20*v2)=(.40+1.20)vf

OpenStudy (fortytherapper):

All that is right. We know v2 though: collides with Glider B ~~at rest~~~

OpenStudy (madinaxxx):

so v2 would be 0?

OpenStudy (fortytherapper):

Yep

OpenStudy (madinaxxx):

(.40*10)+(1.20*0)=(.40+1.20)vf 4+0=(1.60)vf 4=1.6vf 2.5=vf

OpenStudy (fortytherapper):

Looks good to me

OpenStudy (madinaxxx):

thank you!!)))

OpenStudy (fortytherapper):

Anytime